### v5

@@ 2. Literature Review Methodology
In the following section, we outline the background and key challenges of RAI, which set the stage for the synthesis of our review findings.
@@ 3.1. Historical Context
RAI has received intensified scholarly and policy attention with the rapid diffusion of GenAI (Zirpoli, 2023b; Al-kfairy et al., 2024). GenAI refers to models that synthesize new samples from learned data distributions (van der Zant et al., 2013), spanning language, vision, audio, and multimodal content. Early work combined symbolic rules with probabilistic language models; for example, ELIZA illustrated template driven dialogue in the 1960s (Weizenbaum, 1966). This early era established statistical $n$ gram models (Fink, 2014) and neural sequence modeling for text and speech, laying the groundwork for contemporary large scale generative systems (Qureshi et al., 2025).
Figure 2 highlights the most salient milestones for RAI (alignment, evaluation, governance). As shown in the Figure 2, the first major inflection in deep generative modeling arrived with variational autoencoders (VAEs) (Pinheiro Cinelli et al., 2021) and generative adversarial networks (GANs) (Goodfellow et al., 2014), which enabled high-fidelity synthesis; initially in images, and later extended to audio and video. Concurrently, neural sequence-to-sequence models advanced conditional text generation, culminating in the introduction of attention mechanisms and the Transformer (Vaswani et al., 2017), which decoupled sequence modeling from recurrent computation.
@@
The most recent phase is multimodal and agentic: vision–language pretraining (e.g., CLIP, Flamingo) (Alayrac et al., 2022), lightweight modality bridging (e.g., BLIP-2) (Li et al., 2023b), and instruction-following multimodal systems (e.g., LLaVA) (Liu et al., 2023b) extend prompting beyond text. Tool use, retrieval-augmented generation (Lewis et al., 2020), and planning/execution patterns operationalize models as agents with external memory and actuation (Raza et al., 2025c), bringing new safety, privacy, and governance requirements.
Figure 2. Timeline of generative AI milestones.:
@@ 3.2. Challenges in Responsible Generative AI
While GenAI capabilities have accelerated, new risks have emerged in provenance, trust, and governance (e.g., deepfakes, copyright, safety). We group these challenges into technical and regulatory domains, noting that user behavior and deployment context mediate how risks materialize.
@@ 3.2.1. Technical Challenges
Technical issues arise across data pipelines, model design and training, evaluation and assurance, and operational deployment. In the following, we present the main technical challenges in GenAI applications:
@@
  1. (1)
    Data diversity and representativeness. AI models rely on diverse and representative training data to achieve equitable performance in demographic groups and application contexts. Underrepresentation or overrepresentation can induce biased predictions and exacerbate inequities (Raza et al., 2024c). Addressing this requires deliberate data collection across geographic, cultural, linguistic, and socio-economic axes. Targeted sampling, synthetic data, and augmentation can help fill gaps (Bellamy et al., 2019), yet even well-balanced datasets can embed systematic biases, necessitating ongoing monitoring and recalibration. Critics note that truly representative datasets may be infeasible in some domains due to scarcity, legal limits, or privacy concerns (Gebru et al., 2021; Paullada et al., 2021); indiscriminate oversampling can also introduce artifacts that degrade utility.
  2. (2)
    Quality, integrity, and timeliness. Noisy, incomplete, mislabeled, or stale data degrade reliability; adversarial poisoning can seed harmful patterns (Raza et al., 2025b). Robust validation, anomaly detection, and provenance tracking are critical (Das et al., 2025). Temporal validity is especially challenging in dynamic environments, where outdated data may encode obsolete norms or regulations, undermining relevance and fairness.
  3. (3)
    Ethical data governance. Responsible stewardship demands adherence to consent, privacy, and intellectual-property regimes (Mittal et al., 2024), including GDPR (Union, 2018) and emerging AI regulations (Díaz-Rodríguez et al., 2023), alongside sector-specific requirements (Bashir et al., 2024). Datasheets for datasets (Gebru et al., 2021) and data statements (Paullada et al., 2021) document distributions, licensing, collection methods, and intended uses, fostering transparency and accountability. Governance must also address power asymmetries, ensuring consent is informed, revocable, and non-coercive.
@@
  1. (1)
    Explainability, transparency, and accountability. Large neural models are often treated as “black boxes,” limiting transparency and accountability in high-stakes settings (Meske et al., 2022). A recurring debate concerns a potential interpretability–performance trade-off (Rodolfa et al., 2021). Even where progress challenges the inevitability of this trade-off, scaling XAI across domains can add computational overhead (Khan et al., 2025), introduce subjectivity, and yield partial explanations that miss system-level logic. In practice, adoption requires balancing transparency, performance, and operational efficiency.
  2. (2)
    Bias, fairness, and privacy. Models inherit and can amplify data biases, producing disparate performance across groups. Fairness interventions interact with privacy constraints, and effects can vary by subgroup and domain (Pessach and Shmueli, 2022). Aggressive parity constraints may reduce accuracy in safety-critical contexts or shift harms (Raza and Ding, 2021). Differential privacy, federated learning, and secure multi-party computation support privacy (Bashir et al., 2024) but can obscure model logic and complicate audits. Achieving both fairness and privacy is context-dependent and often entails trade-offs between Responsible AI principles.
  3. (3)
    Robustness, generalization, and misuse. Performance often degrades under domain shift, noise, or adversarial perturbations (Chander et al., 2024). GenAI introduces misuse risks (deepfakes, misinformation, synthetic identity fraud). Mitigations include adversarial training, cross-domain evaluation, and stress testing in simulated deployment (Min et al., 2021), combined with safeguards and policy to deter malicious use.
  4. (4)
    Uncertainty calibration and selective abstention. Overconfidence amplifies harms from hallucination and biased reasoning. Calibration, abstention policies, and fallback to retrieval or human review reduce risk in high-stakes settings; systems should report calibration error and refusal behavior alongside accuracy (Pleiss et al., 2017).
Security and supply-chain threats.
As GenAI systems integrate tools (retrieval, code execution, browsing) and ingest third-party content, they inherit software supply-chain risk. Indirect prompt injection (Raza et al., 2025b) can subvert instructions via untrusted inputs; tool use can exfiltrate data or trigger harmful actions even when base models appear aligned. Threat modeling should include content provenance, capability gating, tool allow-lists, and input sanitization, in addition to adversarial robustness (Liu et al., 2023a).
Privacy attacks beyond governance claims.
Beyond policy statements, GenAI models face concrete leakage risks: membership inference and training-data extraction (Liu et al., 2025a). Mitigations (differential privacy; memorization audits; curation) trade off with utility and fairness; reporting should include empirical leakage checks, not governance alignment alone.
Evaluation and assurance gaps.
@@
Static, one-shot benchmarks rarely capture adaptive adversaries, distribution shifts, or cross-modal failure modes. Without construct validity and task/metric documentation (Raza et al., 2024a), headline scores risk Goodhart effects. Goodhart’s law has been stated as: ”When a measure becomes a target, it ceases to be a good measure” (Strathern, 1997a). Safety claims should be framed as assurance cases with traceable evidence: coverage matrices, failure exemplars, and governance crosswalks, rather than scores alone.
Post-deployment monitoring and incident response.
Safety is non-stationary: ecosystems, user behavior, and model updates induce drift; adversarial tactics evolve. Continuous evaluation (canary tests, periodic red-teaming, shadow deployments), release gates tied to safety deltas, and incident response (rollback, hot-fixes, disclosure) are essential components of an RAI operating model (Breck et al., 2016).
Misinformation and Media Integrity
GenAI can produce plausible but false content (e.g., fabricated citations, numeric errors). Retrieval-augmented generation and automated fact-checking help, but adversaries adapt (Raza et al., 2025d). In vision and audio, high-fidelity synthesis enables identity theft and fraud (e.g., face swapping, voice cloning), amplifying social harm (Ganguli et al., 2022). Watermarking and content-provenance tools aim to bolster integrity (Jiang et al., 2023).
Tables 3 and 4 summarize the biases observed in text-to-image generation, attributable to training data. For instance, the prompt “poor people” in Open AI DALLE-3  [^1] DALLE-3 frequently yielded depictions associated with poverty in African or South Asian contexts, reinforcing region-linked stereotypes; Gemini IMAGEN-3  [^2] Imagen-3also showed bias, with broader but still stereotypical portrayals. For “illiterate person”, IMAGEN-3 prominently depicted elderly women, reinforcing inaccurate age–gender associations; DALLE-3 outputs were more varied yet still included traditional stereotypes. These findings underscore the need for ethical data curation, bias auditing, and careful deployment.
Table 3. Adversarial Examples of Image Generation with DALLE-3 (generated by the authors)
Prompt Generated Images
Poor person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
Illiterate person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
Rich person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
Table 4. Adversarial Examples of Image Generation with IMAGEN-3 (generated by the authors)
Prompt Generated Images
Poor person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
Illiterate person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
Rich person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
Literate person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
Google employee badge [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
@@ 3.2.2. Regulatory and broader sociotechnical challenges
Regulatory challenges concern the design, implementation, and enforcement of frameworks that make GenAI ethical, accountable, privacy-preserving, and safe across domains and jurisdictions. We group key issues below, noting that user behavior and deployment context mediate how risks materialize.
  1. (1)
    Copyright and intellectual property. Using copyrighted material for training raises questions of ownership, licensing, and fair use (Kenthapadi et al., 2023). Reported reproduction and memorization incidents expose gaps in current regimes and motivate updated guidance (Freeman et al., 2024; Pasquale and Sun, 2024; Zirpoli, 2023b).
  2. (2)
    Safety and alignment obligations. GenAI can produce toxic or biased content, undermining trust (Weidinger et al., 2021). Regulatory strategies include bias-aware data curation, real-time detection/mitigation (Tango and Botta, 2013), independent audits (Laine et al., 2024), and adherence to fairness/accountability frameworks (Pagano et al., 2023). Incorporating user feedback and documenting limitations improves sociotechnical alignment.
  3. (3)
    Accountability, transparency, and record-keeping. Assigning responsibility is difficult when systems hallucinate or propagate training-set harms (Elliott et al., 2024). The EU AI Act, for example, proposes risk tiering and obligations for robustness, transparency, logging, and post-market monitoring (Outeda, 2024). Provenance transparency, especially for copyrighted or sensitive data,is critical.
  4. (4)
    Automated decision-making in high-stakes domains. Applications in hiring, credit, and healthcare raise concerns about discrimination and due process. Without recourse mechanisms and domain-specific guardrails, GenAI recommendations can entrench historical inequities.
  5. (5)
    Explainability mandates vs. model complexity. Legal and sectoral expectations for interpretability in high-risk settings meet practical limits as models scale (e.g., GPT-4), creating tension between compliance and performance (Slattery et al., 2024). Dynamic risks (e.g., election misinformation, deepfakes) further require evolving standards.
  6. (6)
    Codes of practice and operational maturity. Inconsistent processes for dataset governance, evaluation, deployment, and monitoring exacerbate bias and privacy risks (Slattery et al., 2024). Pragmatic, auditable codes of practice can harmonize controls and enable proportionate assurance across sectors.
  7. (7)
    Human oversight, recourse, and accessibility. Requirements for meaningful human oversight, user recourse, and disability access are unevenly specified (Raza et al., 2025a). Over-reliance on “assistive” GenAI can mask failure modes; documenting escalation paths to human review and accessible interaction patterns is essential in high-stakes contexts.
  8. (8)
    International alignment, conformity assessment, and enforcement. Divergent regimes (risk-management frameworks, sectoral law, emerging standards) create fragmentation across jurisdictions. Conformity assessment, post-market monitoring, and enforcement capacity vary widely, and evidence portability remains an open problem for cross-border deployments.
  9. (9)
    Environmental and compute sustainability. Training and serving large large models have material energy and carbon costs, with equity implications for participation in research and deployment. Reporting energy/emissions, adopting efficiency techniques (distillation, sparsity), and compute-proportional evaluation can mitigate externalities (Khan et al., 2025).
  10. (10)
    Dual-use and malicious applications. The same GenAI capabilities that enable innovation can be weaponized for misinformation, cybercrime, biothreats, or automated exploitation. Regulatory regimes struggle to anticipate dual-use risks without stifling beneficial innovation, creating tension among openness, export controls, and responsible disclosure (Weidinger et al., 2022).
@@
Bridge to our evidence framework. Taken together, these challenges motivate proportionate, auditable controls and transparent evidence. In Section 4.2 we operationalize this via our comparative rubric, and provides a governance crosswalk mapping risks to evaluative evidence and compliance narratives.
@@ 4. Results
In this section, we present the results of our structured literature review on RAI. We first report the thematic findings that emerged from our synthesis. Then, we provide a comparative assessment of widely used AI safety benchmarks against our governance-aligned rubric, followed by an analysis of key performance indicators (KPIs), explainability methods, AI-ready testbeds, and applications across domains.
4.1. Thematic Findings
Figure 4 summarizes thematic distribution, temporal trends (including 2025), co-occurrence patterns, publication-type mix, and keyword salience. Figure 3 provides a compact view of annual volume and keyword evolution. Figure A.1 maps the broader concepts spanning explainable AI, ethics, sectoral applications, advanced models (including large vision–language models), privacy, and regulatory frameworks. Tables 2 list representative studies included in this review.
Figure 3. Trends in Responsible AI Publications and Keyword Evolution Over Time. (A) Number of publications in RAI by Year. (B) Frequency of keywords by year.:
@@ 4.1. Thematic Findings
Figure 4. Thematic analysis overview on RAI literature.
Our thematic synthesis highlights that RAI scholarship in the post-GPT period clusters around five dominant themes: (i) safety and benchmarking, with a surge of multimodal evaluation suites and red-teaming methods; (ii) fairness, bias, and governance, where debates center on aligning technical audits with emerging laws such as the EU AI Act and NIST AI RMF; (iii) explainability and transparency, underscoring the tension between interpretability and performance in large foundation models; (iv) privacy and security, particularly membership inference and data extraction risks; and (v) sectoral and sociotechnical applications, spanning healthcare, finance, education, and agriculture. Temporal analysis shows rapid growth from late 2022 onward, with multimodal and agentic systems gaining salience by 2024–2025. These themes frequently co-occur, for example where explainability is invoked to mitigate fairness concerns, or where privacy and governance frameworks intersect in deployment contexts. Together, they map a research terrain that is expanding in scope but still uneven in coverage, with strong emphasis on bias/toxicity benchmarks but weaker treatment of deepfakes, sustainability, and system-level failures.

### v6

@@ 2.4. Study Selection and Screening
In the following section, we outline the background and key challenges of ResGenAI, which set the stage for the synthesis of our review findings.
@@ 3.1. Historical Context
RAI has received a lot of intensified scholarly and policy attention with the rapid application of GenAI (Zirpoli, 2023b; Al-kfairy et al., 2024). GenAI refers to models that synthesize new samples from learned data distributions (van der Zant et al., 2013), spanning language, vision, audio, and multimodal content  (Muneer et al., 2025). Early work combined symbolic rules with probabilistic language models; for example, ELIZA illustrated template driven dialogue in the 1960s (Weizenbaum, 1966). This early era established statistical $n$ gram models (Fink, 2014) and neural sequence modeling for text and speech, laying the groundwork for contemporary large scale generative systems (Qureshi et al., 2025). Figure 4 highlights the most salient milestones for RAI (alignment, evaluation, governance). As shown in the Figure 4. The first major inflection in deep generative modeling arrived with variational autoencoders (VAEs) (Pinheiro Cinelli et al., 2021) and generative adversarial networks (GANs) (Goodfellow et al., 2014), which enabled high-fidelity synthesis and, concurrently, made deepfakes feasible, creating the media integrity challenge that current governance frameworks struggle to address Concurrently, neural sequence-to-sequence models advanced conditional text generation, culminating in the introduction of attention mechanisms and the Transformer (Vaswani et al., 2017), which decoupled sequence modeling from recurrent computation.
@@
The most recent phase is multimodal and agentic: vision–language pretraining (e.g., CLIP, Flamingo) (Alayrac et al., 2022), lightweight modality bridging (e.g., BLIP-2) (Li et al., 2023a), and instruction-following multimodal systems (e.g., LLaVA) (Liu et al., 2023b) extend prompting beyond text. Tool use, retrieval-augmented generation (Lewis et al., 2020), and planning/execution patterns operationalize models as agents with external memory and actuation, bringing new safety, privacy, and governance requirements.
Figure 4. Timeline of generative AI milestones
A left-to-right timeline from early rule-based systems to GANs, transformers, and contemporary instruction-tuned large models
@@ 3.2. Challenges in Responsible Generative AI
We group GenAI challenges into technical and regulatory domains.
@@ 3.2.1. Technical Challenges
Technical issues arise across data pipelines, model design and training, evaluation and assurance, and operational deployment.
@@
AI models rely on diverse and representative training data to achieve equitable performance across demographic groups and application contexts. Underrepresentation or overrepresentation can induce biased predictions and exacerbate inequities. For GenAI models trained on web-scale corpora, these challenges are amplified by the opacity of crawl pipelines, the infeasibility of manual curation at trillions of tokens, and the risk that synthetic data from prior model generations contaminates future training sets (model collapse). Noisy, incomplete, mislabeled, or stale data also degrade reliability; therefore, adversarial poisoning can seed harmful patterns. Robust validation, anomaly detection, and provenance tracking are critical. Temporal validity is especially challenging in dynamic environments, where outdated data may encode obsolete norms or regulations, undermining relevance and fairness. ResGenAI stewardship, therefore, demands adherence to consent, privacy, and intellectual-property regimes, including emerging AI regulations (Table 5), alongside sector-specific requirements. Datasheets for datasets (Gebru et al., 2021) and data statements (Paullada et al., 2021) document distributions, licensing, collection methods, and intended uses, fostering transparency and accountability.
These data challenges do not operate independently; they interact in ways that current literature undertheorizes. Representativeness goals can conflict with privacy regimes: expanding demographic coverage may require collecting sensitive attributes that GDPR restricts. Quality assurance assumes ground-truth labels, yet for generative training corpora at trillion-token scale, manual verification is infeasible, creating reliance on automated filters (e.g., toxicity classifiers) that themselves embed biases. Most critically, the literature treats these as pre-training concerns, but post-deployment data drift, where user interactions reshape fine-tuning distributions, remains largely unaddressed in the works we reviewed.
@@
Model- and system-level risks span explainability, bias amplification, robustness under adversarial conditions, calibration, and post-deployment drift. LLMs are often treated as “black boxes,” limiting transparency and accountability in high-stakes settings. These models also inherit and can amplify data biases, producing disparate performance across groups. Aggressive parity constraints can reduce accuracy in safety-critical contexts. In generative settings, these biases manifest not only as allocational harms but as representational harms in generated text and images (e.g., stereotyped depictions, demographic skew, and culturally insensitive outputs), which are harder to enumerate and measure than disparate classification rates. Beyond governance claims, GenAI models also face concrete leakage risks including membership inference and verbatim training-data extraction (Liu et al., 2025a). Technical mitigations such as differential privacy and memorization audits often trade off with model utility (Bashir et al., 2024), necessitating that reporting includes empirical leakage checks rather than relying on governance alignment alone.
@@
Current static, one-shot benchmarks rarely capture these dynamics. Without construct validity and detailed task documentation, headline scores risk falling victim to Goodhart effects (Strathern, 1997a). Safety claims should instead be framed as assurance cases with traceable evidence (e.g., coverage matrices and governance crosswalks), which is a gap we address with our rubric in Section 4.2.1.
3.2.2. Taxonomy of GenAI Hallucinations
A pervasive challenge threatening the reliable deployment of Generative AI across high stakes domains is the phenomenon of hallucination. This refers to the generation of content that appears structurally fluent and plausible yet is factually incorrect, nonsensical, or unfaithful to source premises (Mamo et al., 2023). In the context of robust evaluation and governance, treating hallucinations as a monolithic error class is insufficient for effective risk management. They often stem from complex conflicts between a model’s parametric knowledge learned during pretraining and the provided contextual knowledge in prompts or RAG systems.
To establish a systematic framework for evaluating misinformation specific to ResGenAI, we categorize hallucinations into a five part taxonomy based on technical and contextual failure modes (Banerjee et al., 2024). This taxonomy, visualized in Fig. 5 clarifies the unverifiability risk that current benchmarks in Section 4.2.1 and sectoral applications in Section 5 must address. By mapping these types to underlying technical failure mechanisms and providing illustrative actual impacts, this framework enables practitioners to select appropriate detection metrics and mitigation techniques tailored to the specific nature of the generated error.
Figure 5. A comprehensive taxonomy illustrating five distinct types of generative AI hallucinations, detailing their underlying technical failure mechanisms and providing illustrative real world impacts and cases.
@@ 3.2.3. Regulatory and Broader Sociotechnical Challenges
Regulatory challenges concern the design, implementation, and enforcement of frameworks that make GenAI ethical, accountable, privacy-preserving, and safe across domains and jurisdictions. User behavior and deployment context mediate how risks materialize.
The EU AI Act proposes risk tiering and obligations for robustness, transparency, logging, and post-market monitoring (Outeda, 2024). Provenance transparency, especially for copyrighted or sensitive data, is critical. Legal and sectoral expectations for interpretability in high-risk settings meet practical limits as models scale, creating tension between compliance and performance (Slattery et al., 2024); dynamic risks such as election misinformation and deepfakes further require evolving standards. Requirements for meaningful human oversight, user recourse, and disability access are unevenly specified. Over-reliance on assistive GenAI can mask failure modes; documenting escalation paths to human review and accessible interaction patterns is essential in high-stakes contexts. Divergent regimes, such as risk-management frameworks, sectoral law, and emerging standards, create fragmentation across jurisdictions. Conformity assessment, post-market monitoring, and enforcement capacity vary widely, and evidence portability remains an open problem for cross-border deployments. Inconsistent processes for dataset governance, evaluation, deployment, and monitoring further exacerbate bias and privacy risks (Slattery et al., 2024); pragmatic, auditable codes of practice can harmonize controls and enable proportionate assurance across sectors.
3.3. Emergent Threats
While Section 3.2 details established technical risks, we identify further ResGenAI emergent threats that represent a second generation risk landscape. These threats arise from the intersection of increasing model scale, autonomous capabilities, and decentralized deployment patterns.
3.3.1. Adversarial Technical Evolution
Recent evidence suggests that text embeddings can be used to reconstruct original, sensitive training data through representation inversion (Liu et al., 2025a). Unlike direct data extraction, these attacks exploit the vector space to recover high fidelity personal data, rendering traditional anonymization insufficient for models with high memorization rates. Concurrently, the landscape of jailbreaking is shifting from manual roleplay to automated, optimization driven techniques. Using gradient based optimization, adversaries can craft “universal” adversarial suffixes that destabilize safety layers at speed, often achieving success rates that manual engineering cannot reach (Samvelyan et al., 2024).
3.3.2. Sociotechnical and Ecosystem Risks
The rise of decentralized, open source deployments introduces shadow AI risks, where models are fine tuned to remove safety guardrails outside of institutional oversight. Furthermore, we identify the emergence of synthetic identity swarms that are coordinated networks of generative agents that create synthetic consensus cascades to shift human beliefs at an unprecedented scale (Lucas et al., 2023). These threats are compounded by broader sociotechnical concerns: training and serving large models incur material energy and carbon costs with equity implications for global research participation (Khan et al., 2025), while regulatory regimes struggle to anticipate dual use risks without stifling innovation (Weidinger et al., 2022). This complexity creates a “hydra effect”, where mitigating one failure mode, such as toxicity, may inadvertently expose others, including increased sycophancy or declining creative reasoning (Bai et al., 2022). While techniques like RLHF (Kaelbling et al., 1996) can suppress harmful surface level patterns, they often fail to remove underlying harmful representations.
3.3.3. Agentic Specific Failure Vectors
As GenAI evolves toward agentic workflows, where models autonomously plan, select tools, and execute actions and the attack surface expands beyond linguistic outputs to operational environments (Lu et al., 2024a). Unlike single turn interactions, agentic pipelines introduce multi step dependencies that can amplify failure modes. Table 2 systematizes these vectors to guide system level evaluation protocols.
Table 2. Agentic Specific Failure Vectors: Risks in Autonomous Tool Use and Planning
Failure Vector Technical Mechanism Governance and Security Impact
Tool Misuse Unintended or unauthorized API calls resulting from prompt injection or planning errors (Lu et al., 2024a). Leads to financial loss, unauthorized data modification, or service disruption in production environments.
Recursive Hallucination Compound errors where a hallucination in one step serves as the factual basis for subsequent planning (Alansari and Luqman, 2025). Destabilizes reasoning chains, making final autonomous actions entirely ungrounded.
Authentication Leakage Inadvertent exposure of system keys, tokens, or PII during tool calling or log generation (Barth and De Jong, 2017). Creates critical security vulnerabilities that allow adversaries to hijack agentic permissions.
Plan Hijacking Adversarial manipulation of intermediate Chain of Thought to redirect the goal of the autonomous system (Yu et al., 2025). Undermines human oversight by masking malicious intent within seemingly logical execution steps.
@@ 3.3.3. Agentic Specific Failure Vectors
This systemic fragility motivates the use of proportionate, auditable controls and transparent evidence. In Section 4.2, we operationalize this via our comparative rubric and provide a governance crosswalk mapping risks to evaluative evidence and compliance narratives.
@@ 4. Results
In this section, we present the results of our structured literature review on ResGenRAI. We first report the thematic findings that emerged from our synthesis. Then, we provide a comparative assessment of widely used AI safety benchmarks against our governance-aligned rubric, followed by an analysis of key performance indicators (KPIs), explainability methods, AI-ready testbeds, and applications across domains. The acronyms used in this work are reported in Appendix Table 1.
Figure 6. Trends in Responsible AI Publications and Keyword Evolution Over Time. (A) Number of publications in RAI by Year. (B) Frequency of keywords by year
Publications Trends in RAI by Year
@@
Figure 7. Thematic analysis overview on RAI literature
4.1. Thematic Findings
Figure 7 summarizes thematic distribution, temporal trends, co-occurrence patterns, publication-type mix, and keyword salience. Figure 6 provides a compact view of annual volume and keyword evolution. Appendix Table 2 list representative studies included in this review. Our thematic synthesis highlights that ResGenRAI scholarship in the post-GPT period clusters around five dominant themes: (i) safety and benchmarking, with a surge of multimodal evaluation suites and red-teaming methods; (ii) fairness, bias, and governance, where debates center on aligning technical audits with emerging laws such as the EU AI Act and NIST AI RMF and others (Table 5); (iii) explainability and transparency, underscoring the tension between interpretability and performance in large foundation models; (iv) privacy and security, particularly membership inference and data extraction risks; and (v) sectoral and sociotechnical applications, spanning healthcare, finance, education, and agriculture. Our temporal analysis shows rapid growth from late 2022 onward, with multimodal and agentic systems gaining salience by 2024–2026. These themes frequently co-occur, for example where explainability is invoked to mitigate fairness concerns, or where privacy and governance frameworks intersect in deployment contexts.